OpenStudy (anonymous):

The current risk-free rate of return, rRF, is 4 percent and the market risk premium, RPM, is 5 percent. If the beta coefficient associated with a firm’s stock is 2.0, what should be the stock’s required rate of return?

OpenStudy (anonymous):

re = rf + B(rp) re = 4 + 2(5) re = 14%